# A generalization of the relaxation-based matrix splitting iterative method for solving the system of generalized absolute value equations

### v1

@@ 4 Numerical results
In this section, we will use two examples to numerically illustrate the performance of our method for solving the GAVEs (1.1). All experiments are supported by a personal computer with $3.20$ GHZ central processing unit (Intel (R), Corel(TM), i5555-$11320$H), $16$ GB memory and windows $11$ operating system, and implemented in MATLAB $9.10.0$ (R $2021$a). Nine iterative schemes will be tested.
@@
    GRMS: The proposed iterative scheme (2.3) with $M=D-\dfrac{3}{4}L$, $Q=10.5I$ and $Q_{1}=10I$. In our experiments, we let $H=-\theta\zeta I$ with $\zeta\in\mathbb{R}$ being a constant.
@@
    Picard: The Picard iterative scheme (1.5).
@@
    SOR: The SOR-like iterative scheme (2.10).
@@
    MFPI: The modified fixed point iterative scheme (2.11) with $Q_{\rm MFPI}=Q$.
@@
    SSMN: The shift splitting Newton-type iterative scheme (1.6) with $\Omega_{\rm SSMN}=\text{diag}(A)$.
@@
    MGSOR: The modified generalized SOR-like iterative scheme (2.9) with $Q_{\rm MGSOR}=Q$.
@@
    MAMS: The momentum acceleration-based matrix splitting iterative scheme (2.14) with $M_{\rm MAMS}=M$, $\Omega_{\rm MAMS}=\text{diag}(A)$ and the initial vectors $x^{(1)}=x^{(0)}$.
@@
    NSOR: The new SOR-like iterative scheme (2.13).
@@
Example 4.1.
Consider the GAVEs (1.1) with $A=\tilde{A}+\dfrac{1}{5}I$, where
@@
Numerical results for this example are reported in Table 1, in which $\theta_{opt},\omega_{\rm opt}$, $\tau_{\rm opt}$, $\beta_{\rm opt}$, $\alpha_{\rm opt}$, $\sigma_{opt}$ and $\zeta_{\rm opt}$ are numerically optimal iteration parameters of the tested methods, respectively [^2] The same goes to Exmaple 4.2.. For this example, we do not report the results of the NSOR iteration since it is not convergent whenever $m>2$. It follows from Table 1 that the GRMS iteration is superior to all of the other tested methods in terms of IT and CPU.
Table 1: Numerical results for Example 4.1.
@@
Example 4.2.
Consider the GAVEs (1.1) with the same $A$ in Example 4.1 and
@@
Table 2: Numerical results for Example 4.2.
@@ 5 Conclusions
A generalization of the relaxed-based splitting (GRMS) iteration method for solving the generalized absolute value equations is developed. The GRMS method is a general framework which includes many existing methods as its special cases. The convergence of the proposed method is analysed and numerical results demonstrate the effectiveness of the proposed method. However, according to (iii) and (iv) in Section 2, can we develop a method that contains the MSOR and the MAMS for solving the GAVEs (1.1) with $B$ being singular needs further study.

### v2

@@ 5 Numerical results
In this section, we will use two examples to numerically illustrate the performance of our method for solving GAVEs (1). All experiments are supported by a personal computer with $3.20$ GHZ central processing unit (Intel (R), Corel(TM), i5555-$11320$H), $16$ GB memory and windows $11$ operating system, and implemented in MATLAB $9.10.0$ (R $2021$a). Nine iterative schemes will be tested.
  1. []
@@
    GRMS: The proposed iterative scheme (13) with $M=D-\dfrac{3}{4}L$, $Q=10.5I$ and $Q_{1}=10I$. In our experiments, we let $H=-\theta\zeta I$ with $\zeta\in\mathbb{R}$ being a constant.
@@
    Picard: The Picard iterative scheme (5).
@@
    SOR: The SOR-like iterative scheme (20).
@@
    MFPI: The modified fixed point iterative scheme (21) with $Q_{\rm MFPI}=Q$.
@@
    SSMN: The shift-splitting Newton-type iterative scheme (6) with $\Omega_{\rm SSMN}=\text{diag}(A)$.
@@
    MGSOR: The modified generalized SOR-like iterative scheme (19) with $Q_{\rm MGSOR}=Q$.
@@
    MAMS: The momentum acceleration-based matrix splitting iterative scheme (24) with $M_{\rm MAMS}=M$, $\Omega_{\rm MAMS}=\text{diag}(A)$ and the initial vectors $x^{(1)}=x^{(0)}$.
@@
    NSOR: The new SOR-like iterative scheme (23).
@@
Example 4.
Consider GAVEs (1) with $A=\tilde{A}+\dfrac{1}{5}I$, where
@@
Numerical results for this example are reported in Table 1, in which $\theta_{opt},\omega_{\rm opt}$, $\tau_{\rm opt}$, $\beta_{\rm opt}$, $\alpha_{\rm opt}$, $\sigma_{opt}$ and $\zeta_{\rm opt}$ are numerically optimal iteration parameters of the tested methods, respectively [^1] The same goes to Example 5.. For this example, we do not report the results of the NSOR iteration since it is not convergent whenever $m>2$. It follows from Table 1 that the GRMS iteration is superior to all of the other tested methods in terms of IT and CPU.
Table 1: Numerical results for Example 4.
@@
Example 5.
Consider GAVEs (1) with the same $A$ in Example 4 and
@@
Table 2: Numerical results for Example 5.
@@ 6 Conclusions
A generalization of the relaxed-based splitting (GRMS) iteration method for solving the generalized absolute value equations is developed. The GRMS method is a general framework which includes many existing methods as its special cases. The convergence of the proposed method is analysed and numerical results demonstrate the effectiveness of the proposed method. In addition, a few comparison theorems between GRMS and some existing methods are established. However, according to (iii) and (iv) in Section 2, can we develop a method that contains the MSOR and the MAMS for solving GAVEs (1) with $B$ being singular needs further study.
Declarations
  • Funding C. Chen was supported partially by the Fujian Alliance of Mathematics (Grant No. 2023SXLMQN03) and the Natural Science Foundation of Fujian Province (Grant No. 2021J01661). D. Han was supported partially by the National Natural Science Foundation of China (12131004, 11625105) and the Ministry of Science and Technology of China (2021YFA1003600).